OpenStudy (anonymous):

48x^2+72x+9 (factor completely)

OpenStudy (anonymous):

So far I have 3(16x^2 + 24x +3)

myininaya (myininaya):

now let's look to see if 16x^2+24x+3 is factorable a*c=16*3 find two factors of a*c that add up to be b b=24 if there are no factors of a*c, then you say that is far as it can be factored over the integers 48 16*3 8*6 2*24 4*12 can you think any two factors of 48 that add up to be 24

myininaya (myininaya):

we are done since this not factorable over the integers your answer is 3(16x^2+24x+3)

OpenStudy (anonymous):

oh, well thank you :)